# Hyperbolic recurrent neural network as the first type of non-Euclidean neural quantum state ansatz

### v3

@@ 5.3.2 Results
This result is again very interesting in the sense that it might be again indicative of the connection between the geometry of the NQS ansatz used and the structure of the Hamiltonian system under study, similar to the 2D TFIM setting above in which we observe a clear trend of outperformance by 1D hyperbolic GRU compared to 1D Euclidean GRU when there is a hierarchy in the Hamiltonian interaction structure, comprising the first and $N^{th}$ nearest neighbor interactions, introduced by the artificial rearrangement of the 1D spin chain to mimic the 2D spin lattice. In this case, when second nearest neighbor interactions are switched on ($J_{2}>0$), there is also a hierarchy going from the first nearest to the second nearest interactions, and it is our hypothesis that it is the decisive factor that leads to the outperformance of the hyperbolic GRU-based NQS over its Euclidean counterpart. It is unclear to us at the moment, although it is within our expectation, whether this would still be the case in two dimenions, i.e. whether hyperbolic GRU NQS carries on outperforming Euclidean GRU NQS in 2D $J_{1}J_{2}$ models with different 2D lattice geometries (such as square, triangle, kagome, etc.). In order to perform the VMC experiments for 2D $J_{1}J_{2}$ systems, one first needs to generalize the 1D hyperbolic GRU part of the NQS to two dimensions, similar to what was done in [9] in generalizing the 1D Euclidean RNN cell to its 2D version in Eq.(55).
@@ 5.4.1 Experiment details
In the first set of VMC experiments, we only used one variant each for either Euclidean (eGRU-50-s50 with $d_{h}=50$) or hyperbolic GRU (hGRU-50-s50 with $d_{h}=50$) and fixed the number of training epochs to 280 in all cases. In the second set of VMC experiments, we explored the performances of hyperbolic GRU variants with $53\leq d_{h}\leq 58$ to choose the best one among them to compare with the Euclidean GRU benchmark, eGRU-60-s50 with $d_{h}=60$, while extending the number of training epochs to 450 for $(J_{2},J_{3})=(0.2,0.5)$ and 500 for $(J_{2},J_{3})=(0.5,0.2)$. Similar to previous cases, in each experiment, hyperbolic GRU can take up to around 5 times longer to complete the training (on the same hardware) than Euclidean GRU with the same $d_{h}$222222while hyperbolic GRU with smaller $d_{h}$ also takes longer to train, up to around 2-3 times, than Euclidean GRU with larger $d_{h}$. All other training details (optimizer settings, saving-of-best-model-only method) remain the same as in the 1D $J_{1}J_{2}$ case and will not be repeated here.
@@ 5.4.2 Results
From the two sets of VMC experiments for the 1D $J_{1}J_{2}J_{3}$ model, we observe that the obtained results showing that hyperbolic GRU almost always outperformed Euclidean GRU are in agreement with the ones in the 1D $J_{1}J_{2}$ case in which hyperbolic GRU outperfomed its Euclidean version in all cases when $J_{2}\neq 0$, i.e. when the second neighbor interaction is nonvanishing and the Hamiltonian exhibits a hierarchy in the interaction structure comprised of the first and second neighbor interactions. In all four $J_{1}J_{2}J_{3}$ settings where there exist an even more pronounced hierarchy in the interaction structure of the Hamiltonian that comprises the first, second and third nearest neighbor interactions, the clear outperformance of the hyperbolic GRU ansatz reinforces the hypothesis that the hyperbolic geometry of this neural network plays a determining role in handling the type of Hamiltonians with hierarchical structures such as the 1D $J_{1}J_{2}$, $J_{1}J_{2}J_{3}$ systems. As already noted in the 1D $J_{1}J_{2}$ case, an interesting task would be to verify whether hyperbolic GRU carries on outperforming Euclidean GRU for the 2D $J_{1}J_{2}J_{3}$ models on different lattice geometries, but this will require the generalization of the 1D hyperbolic GRU network to a 2D version, similar to the generalization of the 1D Euclidean RNN cell to the 2D version in Eq.55 as done in [9].
@@ 6 Concluding remarks
Given these findings, and the established results in natural language processing that hyperbolic RNNs almost always outperform their Euclidean versions when the training data exhibit hierarchical, tree-like structures [30], [35], [33], it is our hypothesis that hyperbolic RNN-based NQS will likely outperform Euclidean RNN-based NQS in all quantum spin systems in which there is a hierarchical structure in the interaction part of the Hamiltonian. Given the fact that the experiments done in this work have only involved small-size systems due to our limited computing resources, it is obvious that more works are needed to either confirm or refute this hypothesis. If true, this would be an exciting new direction opened up by the possibility of better performances delivered hyperbolic and potentially other non-Euclidean NQS in the problem of approximating the ground state energy of quantum many-body Hamiltonians using VMC - a problem that has been studied using exclusively conventional or Euclidean neural networks.
Several interesting generalizations exist given the results of this work. One immediate and specific direction involves the generalization of the one-dimensional hyperbolic GRU NQS in this work to two dimensions, and compare the performance of this 2D hyperbolic GRU against the 2D Euclidean GRU in the settings of 2D TFIM or 2D Heisenberg $J_{1}J_{2}$ and $J_{1}J_{2}J_{3}$ models. Other broader directions involve the exploration of a different type of hyperbolic RNNs, for example one that based on the Lorentz model of hyperbolic space instead of the Poincaré model [32], in various quantum many-body settings. Yet even more general directions involve the explorations of different types of non-Euclidean NQS beyond hyperbolic RNNs. These new types of non-Euclidean NQS could include, for example, the hyperbolic CNN networks [36] or the hyperbolic transformer neural network [32]. We leave these issues to future works.
@@ Appendix A Two-dimensional RNN model
The 2D Euclidean RNN is a new custom construction from [9] with the following defining equation
@@
Figure 10: The process of rearranging or reshaping a 1D chain of length $N^{2}=(N-1)N-N$ into a 2D lattice $N\times N$ introduces additional interactions between site $i$ and site $(i+N)$ when the vertical dimension is ‘formed’. (a) The original 1D chain with length $N^{2}=(N-1)N+N$, with each site in the chain labeled by their 1D position. (b) The same 1D chain of length $N^{2}=(N-1)N+N$ rearranged in a 2D square lattice of size $N\times N$, with each site in the square lattice labeled by their original 1D position.
@@ Appendix B Convergence curves
Figure 11: Convergence curves of the mean energy from the VMC experiments for 1D TFIM with various NN (clockwise from top: $N=20,40,100,80$ spins) with the three types of NQS ansatzes being eRNN-50-s50, eGRU-50-s50, hypGRU-50-s50. Each subfigure contains a zoomed-in plot showing a narrower range of energy values close to the exact DMRG value.
@@
Figure 12: Convergence curves of the mean energy using 2D Euclidean RNN and 1D Euclidean/hyperbolic GRU ansatzes for the 2D TFIM with ($N_{x},N_{y}$) = (5,5), (7,7), (9,9), (8,8) lattices (clockwise from top). Each subfigure contains a zoomed-in plot showing a narrower range of energy values close to the exact DMRG value.
@@ B.3.2 Convergence curves
Figure 14: Convergence curves of the mean energy using Euclidean GRU and hyperbolic GRU from VMC runs of 1D Heisenberg $J_{1}J_{2}$ model with different $J_{2}$ values. These curves were obtained from training with no gradient clipping. Clockwise from top, $J_{2}$ = 0.0, 0.2, 0.8, 0.5 - for $N=50$ spins.
Figure 15: Convergence curves of the mean energy using Euclidean GRU and hyperbolic GRU from VMC runs of 1D Heisenberg $J_{1}J_{2}$ model with different $J_{2}$ values, with gradient clipping during the training process, - clockwise from top, $J_{2}$ = 0.0, 0.2, 0.8, 0.5 - for $N=50$ spins.
@@
Figure 18: Convergence curves of the mean energy using Euclidean GRU and hyperbolic GRU from VMC runs of 1D Heisenberg $J_{1}J_{2}J_{3}$ model with different $(J_{2},J_{3})$ values - clockwise from top, ($J_{2},J_{3}$) = (0.0,0.5), (0.2,0.2), (0.5,0.2), (0.2,0.5) - for $N=30$ spins.
Figure 19: Convergence curves of the mean energy using Euclidean GRU and hyperbolic GRU from VMC runs of 1D Heisenberg $J_{1}J_{2}J_{3}$ model with different $(J_{2},J_{3})$ values - from top to bottom, ($J_{2},J_{3}$) = (0.2,0.5), (0.5,0.2) - for $N=30$ spins. Each subfigure contains a zoomed-in plot showing a narrower range of energy values close to the exact energy line.

### v4

@@ 5.3.2 Results
This result is again very interesting in the sense that it might be again indicative of the connection between the geometry of the NQS ansatz used and the structure of the Hamiltonian system under study, similar to the 2D TFIM setting above in which we observe a clear trend of outperformance by 1D hyperbolic GRU compared to 1D Euclidean GRU when there is a hierarchy in the Hamiltonian interaction structure, comprising the first and $N^{th}$ nearest neighbor interactions, introduced by the artificial rearrangement of the 1D spin chain to mimic the 2D spin lattice. In this case, when second nearest neighbor interactions are switched on ($J_{2}>0$), there is also a hierarchy going from the first nearest to the second nearest interactions, and it is our hypothesis that it is the decisive factor that leads to the outperformance of the hyperbolic GRU-based NQS over its Euclidean counterpart.
It is unclear to us at the moment, although it is within our expectation, whether this would still be the case in two dimenions, i.e. whether hyperbolic GRU NQS carries on outperforming Euclidean GRU NQS in 2D $J_{1}J_{2}$ models with different 2D lattice geometries (such as square, triangle, kagome, etc.). In order to perform the VMC experiments for 2D $J_{1}J_{2}$ systems, one first needs to generalize the 1D hyperbolic GRU part of the NQS to two dimensions, similar to what was done in [9] in generalizing the 1D Euclidean RNN cell to its 2D version in Eq.(63). However, it is a highly nontrivial task to generalize to the 2D hyperbolic GRU, because even in Euclidean space, the mathematical construction of the 2D GRU is much more complex than that of the 1D GRU. In one dimension, the hidden state $\vec{h}_{i}$ of the 1D GRU only depends on the previous state $\vec{h}_{i-1}$, resulting in the use of a single reset gate $\vec{r}$ and a single update gate $\vec{z}$ (see the GRU defining equations (28)-(28)), but in two dimensions, the hidden state $\vec{h}_{ij}$ of the 2D GRU receives updades from three different previous hidden states $\vec{h}_{i-1,j}$, $\vec{h}_{i,j-1}$ and $\vec{h}_{i-1,j-1}$, resulting in the use of three different reset gates $\vec{r}$’s and four different update gates $\vec{z}$’s (see Eq.56 - Eq.60, which are the defining equations of the 2D Euclidean GRU as constructed in [56]). [^25] A concrete construction of the 2D GRU is given in the equations below from the work [56] $\vec{r}_{l}$ $=$ ${\sigma}\left(W_{r_{l}}\vec{q}+\vec{b}_{r_{l}}\right),\hskip 14.22636pt\vec{r}_{t}={\sigma}\left(W_{r_{t}}\vec{q}+\vec{b}_{r_{t}}\right),\hskip 14.22636pt{}\vec{r}_{d}={\sigma}\left(W_{r_{d}}\vec{q}+\vec{b}_{r_{d}}\right),$ (56) $\vec{z}_{i}$ $=$ ${\sigma}\left(W_{z_{i}}\vec{q}+\vec{b}_{z_{i}}\right),\hskip 14.22636pt\vec{z}_{l}={\sigma}\left(W_{z_{l}}\vec{q}+\vec{b}_{z_{l}}\right),$ (57) $\vec{z}_{t}$ $=$ ${\sigma}\left(W_{z_{t}}\vec{q}+\vec{b}_{z_{t}}\right),\hskip 14.22636pt\vec{z}_{d}={\sigma}\left(W_{z_{d}}\vec{q}+\vec{b}_{z_{d}}\right),$ (58) $\tilde{\vec{h}}_{ij}$ $=$ $f\left(W\,\vec{x}_{ij}+U\left[\vec{r}\oplus\left\{\vec{h}_{i,j-1},\vec{h}_{i-1,j},\vec{h}_{i-1,j-1}\right\}\right]+\vec{b}\right)$ (59) $\vec{h}_{ij}$ $=$ $\vec{z}_{l}\oplus\vec{h}_{i,j-1}+\vec{z}_{t}\oplus\vec{h}_{i-1,j}\oplus\vec{z}_{d}\oplus\vec{h}_{i-1,j-1}+\vec{z}_{i}\oplus\tilde{\vec{h}}_{ij},$ (60) where $\vec{q}=[\vec{h}_{i-1,j},\vec{h}_{i,j-1},\vec{h}_{i-1,j-1},\vec{x}_{ij}]^{T}$ is the concatenated vector containing the 3 previous hidden states $\vec{h}_{i-1,j},\vec{h}_{i,j-1},\vec{h}_{i-1,j-1}$ and the input vector $\vec{x}_{ij}$, $\vec{r}_{l,t,d}$ are the three reset gates corresponding to $\vec{h}_{i,j-1}$, $\vec{h}_{i-1,j}$, $\vec{h}_{i-1,j-1}$, $\vec{z}_{l,t,d,i}$ are the four update gates. The $\vec{W}$’s, $\vec{U}$’s and $\vec{b}$’s are the weight matrices and bias vectors. $\sigma$} is the sigmoid function, $f$ is a nonlinear activation function, e.g. $\tanh$. Even if we were to do away with the hidden state $\vec{h}_{i-1,j-1}$ [^26] The hidden state $\vec{h}_{i-1,j-1}$ might not be necessary in the context of 2D spin lattice with only vertical and horizontal nearest-neighbor interactions. In their construction of the 2D RNN, the authors of [9] also did not make use of $\vec{h}_{i-1,j-1}$, see Eq.63., we would still require two reset gates and three update gates for the 2D GRU. As such, with the many more parameters involved in the construction of the 2D Euclidean GRU and the already-complex definitions of various arithmetic operations in hyperbolic space, the hyperbolic 2D GRU (obtained from converting the 2D Euclidean GRU) will necessarily be much more mathematically complex than the 1D hyperbolic GRU. Such a network will be very computationally costly to train, both in terms of time and the hardware resources required.
@@ 5.4.1 Experiment details
In the first set of VMC experiments, we only used one variant each for either Euclidean (eGRU-50-s50 with $d_{h}=50$) or hyperbolic GRU (hGRU-50-s50 with $d_{h}=50$) and fixed the number of training epochs to 280 in all cases. In the second set of VMC experiments, we explored the performances of hyperbolic GRU variants with $53\leq d_{h}\leq 58$ to choose the best one among them to compare with the Euclidean GRU benchmark, eGRU-60-s50 with $d_{h}=60$, while extending the number of training epochs to 450 for $(J_{2},J_{3})=(0.2,0.5)$ and 500 for $(J_{2},J_{3})=(0.5,0.2)$. Similar to previous cases, in each experiment, hyperbolic GRU can take up to around 5 times longer to complete the training (on the same hardware) than Euclidean GRU with the same $d_{h}$272727while hyperbolic GRU with smaller $d_{h}$ also takes longer to train, up to around 2-3 times, than Euclidean GRU with larger $d_{h}$. All other training details (optimizer settings, saving-of-best-model-only method) remain the same as in the 1D $J_{1}J_{2}$ case and will not be repeated here.
@@ 5.4.2 Results
From the two sets of VMC experiments for the 1D $J_{1}J_{2}J_{3}$ model, we observe that the obtained results showing that hyperbolic GRU almost always outperformed Euclidean GRU are in agreement with the ones in the 1D $J_{1}J_{2}$ case in which hyperbolic GRU outperfomed its Euclidean version in all cases when $J_{2}\neq 0$, i.e. when the second neighbor interaction is nonvanishing and the Hamiltonian exhibits a hierarchy in the interaction structure comprised of the first and second neighbor interactions. In all four $J_{1}J_{2}J_{3}$ settings where there exist an even more pronounced hierarchy in the interaction structure of the Hamiltonian that comprises the first, second and third nearest neighbor interactions, the clear outperformance of the hyperbolic GRU ansatz reinforces the hypothesis that the hyperbolic geometry of this neural network plays a determining role in handling the type of Hamiltonians with hierarchical structures such as the 1D $J_{1}J_{2}$, $J_{1}J_{2}J_{3}$ systems. As already noted in the 1D $J_{1}J_{2}$ case (see Section 5.3.2), an interesting task would be to verify whether hyperbolic GRU carries on outperforming Euclidean GRU for the 2D $J_{1}J_{2}J_{3}$ models on different lattice geometries, but this will require the generalization of the 1D hyperbolic GRU network to a 2D version, similar to the generalization of the 1D Euclidean RNN cell to the 2D version in Eq.63 as done in [9]. As also noted in Section 5.3.2, this generalization is a challenging task, given that the complexity of the resulting 2D hyperbolic network will result in training processes that are highly computationally demanding.
@@ 6 Concluding remarks
At this point, it is useful to recall the established results in natural language processing (NLP) and graph embedding that hyperbolic RNNs almost always outperform their Euclidean versions when the training data exhibit hierarchical structures [30], [36], [34]. In these NLP/graph embedding contexts, the fact that hierarchical structures are often tree-like in nature largely explains the superior performances of hyperbolic neural networks, since a key property of hyperbolic space - the exponential expansion of space at large distances - is responsible for its ability to accomodate or embed tree structures with very low distortions. In the context of physical Hamiltonian spin systems, the hierarchical structures are not really tree-like in nature, but nonetheless a clear hierarchy in the interaction part exists in all the systems that hyperbolic NQS was found to have outperformed its Euclidean counterpart. Thus, it is our hypothesis that hyperbolic RNN-based NQS will likely outperform Euclidean RNN-based NQS in all quantum spin systems in which there is a hierarchical structure in the interaction part of the Hamiltonian. Given the fact that the experiments done in this work have only involved small-size systems due to our limited computing resources, it is obvious that more works are needed to either confirm or refute this hypothesis. If true, this would be an exciting new direction opened up by the possibility of better performances delivered hyperbolic and potentially other non-Euclidean NQS in the problem of approximating the ground state energy of quantum many-body Hamiltonians using VMC - a problem that has been studied using exclusively conventional or Euclidean neural networks.
Several interesting generalizations exist given the results of this work.
  • One immediate and specific direction involves the generalization of the one-dimensional hyperbolic GRU NQS in this work to two dimensions, and compare the performance of this 2D hyperbolic GRU against the 2D Euclidean GRU in the settings of 2D TFIM or 2D Heisenberg $J_{1}J_{2}$ and $J_{1}J_{2}J_{3}$ models. The main challenge for this task lies in the fact that the computational cost required to train the 2D Euclidean and hyperbolic GRUs will be much more intensive than the already-intensive computational cost required to train the 1D hyperbolic GRU. This is because given the fact that the 2D Euclidean GRU is more mathematically complex than the 1D GRU, its hyperbolic conversion will necessarily result in an even more complex construction, as previously noted in detail in Section 5.3.2.
[…]
  • Yet even more general directions involve the explorations of different types of non-Euclidean NQS beyond hyperbolic RNNs. These new types of non-Euclidean NQS could include, for example, the hyperbolic CNN networks [37] or the hyperbolic transformer neural network [33].
We leave these issues to future works.
@@ Appendix A Two-dimensional RNN model
The 2D Euclidean RNN is a new custom construction from [9] with the following defining equation [^29] A more general discussion of the multi-dimensional RNN can be found in [55].
@@
Figure 10: The process of rearranging or reshaping a 1D chain of length $N^{2}=(N-1)N+N$ into a 2D lattice $N\times N$ introduces additional interactions between site $i$ and site $(i+N)$ when the vertical dimension is ‘formed’. In the original 1D lattice in the setting of 1D TFIM, there exist only the nearest-neighbor interactions $\langle i,i+1\rangle$. When the 1D spin chain (with site locations labeled bya single index ii)is reshaped into a 2D square lattice (with site locations labeled by $(i^{\prime},j^{\prime})$) in the setting of 2D TFIM, besides the 1D original $\langle i,i+1\rangle$ interaction which corresponds to the 2D nearest neighbor interaction $\langle(i^{\prime},j^{\prime}),(i^{\prime},j^{\prime}+1)\rangle$ along the horizontal axis, an additional type of interactions $\langle i,i+N\rangle$, which corresponds to the 2D vertical nearest neighbor interaction $\langle(i^{\prime},j^{\prime}),(i^{\prime}+1,j^{\prime})\rangle$, is introduced (see Eq.53 for the 1D-2D mapping and the discussion in Section 5.2.2 for more details). (a) The original 1D chain with length $N^{2}=(N-1)N+N$, with each site in the chain labeled by the 1D index $1\leq i\leq(N-1)N+N$. (b) The same 1D chain of length $N^{2}=(N-1)N+N$ rearranged in a 2D square lattice of size $N\times N$, with each site in the square lattice labeled by their original 1D position.
@@ Appendix B Convergence curves
Figure 11: Convergence curves of the mean energy from the VMC experiments for 1D TFIM with various NN (clockwise from top: $N=20,40,100,80$ spins) with the three types of NQS ansatzes being eRNN-50-s50, eGRU-50-s50, hypGRU-50-s50. Each subfigure contains a zoomed-in plot showing a narrower range of energy values close to the exact DMRG value. The vertical dashed lines (whose legend labels contain the letters ‘bm’) indicate the epochs at which the best models were saved. The epochs at which the best models were saved were automatically determined by the training loop such that after convergence is achieved, the best models not only have the lowest mean energy but also with the variance below the tolerance threshold (see the discussion in Section 5 just before Section 5.1). When $N=100$, the best models for eGRU and eRNN were both saved at epoch 120, thus only one vertical dashed line is visible, since these 2 dashed lines corresponding to eGRU and eRNN coincide with each other.
@@
Figure 12: Convergence curves of the mean energy using 2D Euclidean RNN and 1D Euclidean/hyperbolic GRU ansatzes for the 2D TFIM with ($N_{x},N_{y}$) = (5,5), (7,7), (9,9), (8,8) lattices (clockwise from top). Each subfigure contains a zoomed-in plot showing a narrower range of energy values close to the exact DMRG value. The vertical dashed lines (whose legend labels contain the letters ‘bm’) indicate the epochs at which the best models were saved. The epochs at which the best models were saved were automatically determined by the training loop such that after convergence is achieved, the best models not only have the lowest mean energy but also with the variance below the tolerance threshold (see the discussion in Section 5 just before Section 5.1). For $(N_{x},N_{y})=(5,5)$, the number of training epochs of 2d-eRNN was much smaller than those of 1d-eGRU and 1d-hGRU, so the purple dashed line corresponding to the epoch at which the 2d-eRNN best model was saved appear much earlier those corresponding to 1d-eGRU and 1d-hGRU. When $(N_{x},N_{y})=(8,8)$, the best models for 1d-eGRU and 1d-hGRU were both saved the same epoch, thus only one vertical dashed line is visible, since these 2 dashed lines corresponding to 1d-eGRU and 1d-hGRU coincide with each other.
@@ B.3.2 Convergence curves
Figure 14: Convergence curves of the mean energy using Euclidean GRU and hyperbolic GRU from VMC runs of 1D Heisenberg $J_{1}J_{2}$ model with different $J_{2}$ values. These curves were obtained from training with no gradient clipping. Clockwise from top, $J_{2}$ = 0.0, 0.2, 0.8, 0.5 - for $N=50$ spins. The vertical dashed lines (whose legend labels contain the letters ‘bm’) indicate the epochs at which the best models were saved. The epochs at which the best models were saved were automatically determined by the training loop such that after convergence is achieved, the best models not only have the lowest mean energy but also with the variance below the tolerance threshold (see the discussion in Section 5 just before Section 5.1). In this case, there are various kinks or jumps in the curves, indicating instabilities during the training process, espcially those corresponding to the Euclidean network. When $J_{2}=0.2$, the best model was saved right before there was a large jump in the energy.
Figure 15: Convergence curves of the mean energy using Euclidean GRU and hyperbolic GRU from VMC runs of 1D Heisenberg $J_{1}J_{2}$ model with different $J_{2}$ values, with gradient clipping during the training process, - clockwise from top, $J_{2}$ = 0.0, 0.2, 0.8, 0.5 - for $N=50$ spins. The vertical dashed lines (whose legend labels contain the letters ‘bm’) indicate the epochs at which the best models were saved. The epochs at which the best models were saved were automatically determined by the training loop such that after convergence is achieved, the best models not only have the lowest mean energy but also with the variance below the tolerance threshold (see the discussion in Section 5 just before Section 5.1).
@@
Figure 18: Convergence curves of the mean energy using Euclidean GRU and hyperbolic GRU from VMC runs of 1D Heisenberg $J_{1}J_{2}J_{3}$ model with different $(J_{2},J_{3})$ values - clockwise from top, ($J_{2},J_{3}$) = (0.0,0.5), (0.2,0.2), (0.5,0.2), (0.2,0.5) - for $N=30$ spins. The vertical dashed lines (whose legend labels contain the letters ‘bm’) indicate the epochs at which the best models were saved. The epochs at which the best models were saved were automatically determined by the training loop such that after convergence is achieved, the best models not only have the lowest mean energy but also with the variance below the tolerance threshold (see the discussion in Section 5 just before Section 5.1).
Figure 19: Convergence curves of the mean energy using Euclidean GRU and hyperbolic GRU from VMC runs of 1D Heisenberg $J_{1}J_{2}J_{3}$ model with different $(J_{2},J_{3})$ values - from top to bottom, ($J_{2},J_{3}$) = (0.2,0.5), (0.5,0.2) - for $N=30$ spins. Each subfigure contains a zoomed-in plot showing a narrower range of energy values close to the exact energy line. The vertical dashed lines (whose legend labels contain the letters ‘bm’) indicate the epochs at which the best models were saved. The epochs at which the best models were saved were automatically determined by the training loop such that after convergence is achieved, the best models not only have the lowest mean energy but also with the variance below the tolerance threshold (see the discussion in Section 5 just before Section 5.1).